If 22 L of helium at 20.0°C is allowed to expand to 44 L, with the pressure remaining the same, what is the new temperature in Celsius?


1
Expert's answer
2021-03-24T06:41:11-0400

Using Charle's Law,

V1 = 22L

T1 = 20°C + 273 = 293K

V2 = 44L

T2 = ?


"\dfrac{V_1}{T_1} = \dfrac{V_2}{T_2}"


"\dfrac{22L}{293K} = \dfrac{44L}{T_2}"


"T_2 = \dfrac{293×44}{22} = 586K"


T2 (°C) = 586K -273 = 313°C